An inner produce, such as the dot product, always uses two vectors as operands and produces a scalar number as the result. O True O False
Answer #1

True

The dot product v.w of two vectors is also called the inner product or the scalar product.

The vectors are usually written in Hamilton's quaternions. for ex : 3 i + 2 j + 5 k

If you multiply two vectors together,you’ll get the scalar part of the product which is the negation of the dot product.The dot product of two vectors is a scalar because it was the part of the quaternion product that was a scalar.

i * i = -1

j * j = -1

k * k = -1